Hello There,
I need help solving for x and then classifying the triangle by its angles.(I already know how to classify them) My question is: Find the value of X.Then Classify the triangle by its angle measures.
angle#1:(2x+11) angle#2:x angle#3:94
What Ive tried: I set it up in a equation:94+x+(2x+11)=180
I combined like terms (2x+x)and i got 3x. Then i did 94+3x+11=180
Next,I tried 180-2=178...from that point on,..iv'e been stuck a while. And thats only number 1. Thank for helping ^_^

Answer by belle3833(46)

From the point where you have 94+3x+11=180
(94+11)+3x=180
105+3x=180
3x=180-105
3x=75
x=75/3
x=25
angle 1 =2x+11=2 x 25+11=61
angle2=x=25
angle 3=94
Hope this helps